OpenStudy (anonymous):

Bob bought a glass pyramid from a gift store. The pyramid, with a height of 5 cm and a square base of side 6 cm, was packed in a box shaped as a cube of side 7 cm. The empty space in the box was packed with paper tissue. How many cubic centimeters of the box were stuffed with paper tissue? A 283 cubic centimeters B 293 cubic centimeters C 313 cubic centimeters D 343 cubic centimeters

OpenStudy (anonymous):

A. 283 cm^3

OpenStudy (anonymous):

How did you get that though?

OpenStudy (anonymous):

The cube has a volume s^3 =7^3=343. The pyramid had a volume of BH/3 where B is the area of the base and h is the height. B is 6*6=36, h=5 so V=60 Vtp=Vc-Vp Vtp=343-60=283 cm^3

OpenStudy (anonymous):

What about this question? I got the answer 8 times is greater (B) but I keep second guessing it. A floor has two square-shaped designs. The area of the second square-shaped design is sixteen times greater than the area of the first square-shaped design. Which statement gives the correct relationship between the lengths of the sides of the two squares? A The length of the side of the second square is 2 times greater than the length of the side of the first square. B The length of the side of the second square is 8 times greater than the length of the side of the first square. C The length of the side of the second square is 16 times greater than the length of the side of the first square. D The length of the side of the second square is 4 times greater than the length of the side of the first square.

OpenStudy (anonymous):

D. The length of the side of the second square is 4 times greater than the length of the side of the first square.

OpenStudy (anonymous):

Wait for one sec, double checking.....

OpenStudy (anonymous):

I got D. as the answer again.

OpenStudy (anonymous):

ok and how did you get that?

OpenStudy (anonymous):

The ratio of the sides would be the square root of of the ratios of the areas. Square root of 16 = 4. 4*4=16

OpenStudy (anonymous):

I had to redo it but I'm pretty sure that's right.

OpenStudy (anonymous):

ok thanks. I must have done something wrong :)

OpenStudy (anonymous):

No problem. I do that all the time to :P

OpenStudy (anonymous):

Wish you luck on your test. :)